OpenStudy (anonymous):

Solve the system: –5x – y = –11 2x + 2y = –10 What is the value of y?

OpenStudy (anonymous):

Is this the only information given

OpenStudy (tkhunny):

Multiply the second equation by 5/2 and then think about the solution.

OpenStudy (anonymous):

yes that's the question. that's some bull ain't it lol

OpenStudy (tkhunny):

Seriously, 5/2.

OpenStudy (anonymous):

(–5x – y = –11) 2 (2x + 2y = –10)5 -10x-2y=-22 + 10x+10y=-50 [8y=-72] 1/8 y=-9

OpenStudy (anonymous):

thank you bahupepe

OpenStudy (anonymous):

urwlcome